# Optimizing Multi-Stuttered Speech Classification: Leveraging Whisper’s Encoder for Efficient Parameter Reduction in Automated Assessment

### v2

@@ 3.3 Model Architecture
Whisper is an encoder-decoder architecture based on weakly supervised transformers that are state-of-the-art at the moment, particularly for speech recognition [7]. Whisper uses a 16 kHz sampling rate for the audio clips that are fed into it. Next, these audio clips are transformed into a log Mel spectrogram representation. After that, two 1-D convolutional layers are used to process them in order to extract features. Then, depending on the requirements of a given task, these features are fed into an encoder and a decoder. To enable the model to be used for a variety of use cases, such as speech recognition, language identification, voice detection, and speech translation, the decoder is trained in a multitask format. However, for the classification of multi-stuttered disfluency types, only the encoder part of the model is required.
@@ 3.3.8 Loss Function
Loss functions play a significant role in the evaluation of a model as it quantifies between predictions and actual results. Since our problem is a multi-label classification of multiple disfluences, the Binary Cross-Entropy (BCE) loss function is well-suited [25]. More specifically, we have used BCEWithLogitsLoss function in Pytorch [^3] https://pytorch.org/docs/stable/generated/torch.nn.BCEWithLogitsLoss.html, it combines sigmoid function and BCE Loss in a single operation. Therefore, it can be acquired for instances which have multiple labels. The mathematical formula for BCE Loss is as follows
@@ 3.4 Experimental Design
After curating the dataset for multi-stuttered speech, speaker-exclusive data splits discussed in the section 3.1 are then fine-tuned. In this study, to cater for multi-stuttered instances, encoder-only Whisper has been leveraged by fine-tuning its pre-trained English base model. In order to adapt the model to our specific task, we used Hugging Face Library to acquire its module namely, Whisper for Audio Classification.  [^4] https://huggingface.co/docs/transformers/model_doc/Whisper.
@@
First, the base models of Wav2vec2.0 and Whisper were finetuned on the multi-stuttered dataset on all the splits. Based on the results, Whisper was chosen for further experimentation. Next, the encoder layer optimization strategy was applied, in which the first three encoder layers of the Whisper model were frozen while fine-tuning. The insights of the optimization strategy were gained from earlier research [8]. For further reduction of trainable parameters, we have also experimented with the freezing of the feature extractor. In addition, we have also included the experimentation of freezing the first four, first five layers, and all layers of the encoder. Lastly, the Whisper model with only four encoder layers was also fine-tuned. However, the results exhibited a significant decline, which led us to discontinue further experimentation. Detailed outcomes stemming from this experimental design are presented and discussed in the section 4.
@@ 4 Results and Discussion
In this section, we begin by analyzing the outcomes between the Wav2vec2.0 and Whisper models. The results led us to continue with the Whisper model for further experimentation to achieve the classification of multi-stuttered disfluencies in speech. Next, we have explored the impact of freezing the feature extractor on the outcome. The findings demonstrated that freezing the feature extractor does not affect the results, thereby preferring the respective configuration with reduced trainable parameters. Finally, building on the paper [8], we extended our experimentation to the remaining encoder layers. Through these experiments, we identified the optimal layer freezing configuration of the model, and compared with with [10].
In the literature, the research on multi-stuttered classification of disfluency is notably scarce. In the study [10], authors have focused on this problem leveraging Wav2vec2.0 for the classification of disfluency types in multi-stuttered speech. However, in the paper, [8] Whisper appeared to be a more suitable choice for this task. To make a concrete choice between these two models we fine-tuned both Wav2vec2.0 and Whisper. The comparative analysis in table 5 demonstrates that Whisper is a better choice across all the data splits in contrast to Wav2vec2.0. For this experimentation, Wav2vec2.0’s base model which consists of 12 encoder layers was acquired. In contrast, Whisper’s base model consisting of 6 encoder layers was fine-tuned. All the Layers were kept unfrozen in both of the models. It is noticeable in the table 5 that there is a considerable variation of results between the models. Given their architectures, one of the key differences between Wav2vec2.0 and Whisper is their approaches to input representation of waveforms, potentially leading to substantial differences in results. Whisper converts the raw input into mel-spectrograms which are then passed to convolutional layers for feature extraction. On the other hand, Wav2vec2.0 directly extracts these features from raw audio through convolutional layers. Applying the convolutional layers directly on raw waveforms extracts more features as they are learned through trainable parameters. It seems that acquiring features through convolutional should have better outcomes as it is learning more features but it may not be necessary. For the respective task, features that are related to human speech are more relevant and also limit the computational head on the models. Therefore, one of the major reasons Whisper is performing better could be due to the usage of mel-spectrograms for feature extraction. Mel-spectrograms mimic the human auditory system, and these finite features are then passed on to convolutional layers and eventually to the encoder layers.
@@
Note: The terms micro, macro, and weighted” refer to the types of F1-scores.
Since it has been established that Whisper is the better choice for the task at hand. Based on the groundwork laid by [8], it was proved through experimentation that freezing the initial layers of the encoder model will yield better results. Consequently, this methodology will have fewer trainable parameters. Therefore, we experimented with this particular configuration of freezing the first three layers of the model frozen. Building on this study, we also experimented with freezing the feature extractor for further reduction of parameters. The results presented in the table 6 demonstrate that the outcomes of the freezed feature extractor are comparable to the unfreezed feature extractor. Thus, the rest of the experiments are done keeping the feature extractor layers frozen. Table 7 shows the comparison between the reduced learnable parameters.
Table 6: Comparative Analysis of Model Performance with Frozen and Unfrozen Feature Extractor Across SEP28k-E and FluencyBank Datasets
@@
Note: In the table, ”UnFrz0-5” denotes layers 0 to 5 are unfrozen (trainable), ”UnFrz0-5+FrzFE” indicates layers 0 to 5 are unfrozen with the feature extractor (FE) being frozen, ”Frz0-2” signifies only layers 0, 1, and 2 are frozen (not trainable), and ”Frz0-2+FrzFE” means layers 0, 1, and 2 are frozen along with the feature extractor.
@@
We have also compared our results from the top five frozen layers with the outcomes reported by [25]. The figure 4 illustrated that our model outcomes have outperformed the Wav2vec2.0 proposed by [25] with considerably less trainable parameters. In addition, it’s noteworthy that the Whisper model’s variant with 6 encoder layers was used in our experiments, and Wav2vec2.0 model’s large variant which consists of 24 encoder layers was used for the respective task. Therefore, Wav2vec2.0 large variant had 315.70 million trainable parameters, and our proposed whisper model configuration had 20.72 million trainable parameters.

### v3

@@ 3.3 Model Architecture
Whisper is an encoder-decoder architecture based on weakly supervised transformers that are state-of-the-art at the moment, particularly for speech recognition Radford et al., (2023). Whisper uses a 16 kHz sampling rate for the audio clips that are fed into it. Next, these audio clips are transformed into a log Mel spectrogram representation. After that, two 1-D convolutional layers are used to process them in order to extract features. Then, depending on the requirements of a given task, these features are fed into an encoder and a decoder. To enable the model to be used for a variety of use cases, such as speech recognition, language identification, voice detection, and speech translation, the decoder is trained in a multitask format. However, for the classification of multi-stuttered disfluency types, only the encoder part of the model is required.
@@ 3.3.8 Loss Function
Loss functions play a significant role in the evaluation of a model as it quantifies between predictions and actual results. Since our problem is a multi-label classification of multiple disfluences, the Binary Cross-Entropy (BCE) loss function is well-suited Bayerl et al., 2023b . More specifically, we have used BCEWithLogitsLoss function in Pytorch  [^3] https://pytorch.org/docs/stable/generated/torch.nn.BCEWithLogitsLoss.html, it combines sigmoid function and BCE Loss in a single operation. Therefore, it can be acquired for instances which have multiple labels. The mathematical formula for BCE Loss is as follows
@@ 3.4 Experimental Design
After curating the dataset for multi-stuttered speech, speaker-exclusive data splits discussed in the section 3.1 are then fine-tuned. In this study, to cater for multi-stuttered instances, encoder-only Whisper has been leveraged by fine-tuning its pre-trained English base model. In order to adapt the model to our specific task, we used Hugging Face Library to acquire its module namely, Whisper for Audio Classification  [^4] https://huggingface.co/docs/transformers/model_doc/Whisper.
@@
First, the base models of Wav2Vec2.0 and Whisper were finetuned on the multi-stuttered dataset on all the splits. Based on the results, Whisper was chosen for further experimentation. Next, the encoder layer optimization strategy was applied, in which the first three encoder layers of the Whisper model were frozen while fine-tuning. The insights of the optimization strategy were gained from earlier research Ameer et al., (2023). For further reduction of trainable parameters, we have also experimented with the freezing of the feature extractor. In addition, we have also included the experimentation of freezing the first four, first five layers, and all layers of the encoder. Lastly, the Whisper model with only four encoder layers was also fine-tuned. However, the results exhibited a significant decline, which led us to discontinue further experimentation. Detailed outcomes stemming from this experimental design are presented and discussed in the section 4.
@@ 4 Results and Discussion
In this section, we begin by analyzing the outcomes between the Wav2Vec2.0 and Whisper models. The results led us to continue with the Whisper model for further experimentation to achieve the classification of multi-stuttered disfluencies in speech. Next, we have explored the impact of freezing the feature extractor on the outcome. The findings demonstrated that freezing the feature extractor does not affect the results, thereby preferring the respective configuration with reduced trainable parameters. Finally, building on the paper Ameer et al., (2023), we extended our experimentation to the remaining encoder layers. Through these experiments, we identified the optimal layer freezing configuration of the model, and compared with with Bayerl et al., 2023b .
In the literature, the research on multi-stuttered classification of disfluency is notably scarce. In the study Bayerl et al., 2023b , authors have focused on this problem leveraging Wav2Vec2.0 for the classification of disfluency types in multi-stuttered speech. However, in the paper, Ameer et al., (2023) Whisper appeared to be a more suitable choice for this task. To make a concrete choice between these two models we fine-tuned both Wav2Vec2.0 and Whisper. The comparative analysis in Table 5 demonstrates that Whisper is a better choice across all the data splits in contrast to Wav2Vec2.0. For this experimentation, Wav2Vec2.0’s base model which consists of 12 encoder layers was acquired. In contrast, Whisper’s base model consisting of 6 encoder layers was fine-tuned. All the Layers were kept unfrozen in both of the models. It is noticeable in the Table 5 that there is a considerable variation of results between the models. Given their architectures, one of the key differences between Wav2Vec2.0 and Whisper is their approaches to input representation of waveforms, potentially leading to substantial differences in results. Whisper converts the raw input into mel-spectrograms which are then passed to convolutional layers for feature extraction. On the other hand, Wav2Vec2.0 directly extracts these features from raw audio through convolutional layers. Applying the convolutional layers directly on raw waveforms extracts more features as they are learned through trainable parameters. It seems that acquiring features through convolutional layers should have better outcomes as it is learning more features but it may not be necessary. For the respective task, features that are related to human speech are more relevant and also limit the computational head on the models. Therefore, one of the major reasons Whisper is performing better could be due to the usage of mel-spectrograms for feature extraction. Mel-spectrograms mimic the human auditory system, and these finite features are then passed on to convolutional layers and eventually to the encoder layers.
@@
Note: The terms micro, macro, and weighted refer to the types of F1-scores.
Table 6: Comparative Analysis of Model Performance with Frozen and Unfrozen Feature Extractor Across SEP-28k-E and FluencyBank Datasets
@@
Note: In the table, "UnFrz0-5" denotes layers 0 to 5 are unfrozen (trainable), "UnFrz0-5+FrzFE" indicates layers 0 to 5 are unfrozen with the feature extractor (FE) being frozen, "Frz0-2" signifies only layers 0, 1, and 2 are frozen (not trainable), and "Frz0-2+FrzFE" means layers 0, 1, and 2 are frozen along with the feature extractor.
Since it has been established that Whisper is the better choice for the task at hand. Based on the groundwork laid by Ameer et al., (2023), it was proved through experimentation that freezing the initial layers of the encoder model will yield better results. Consequently, this methodology will have fewer trainable parameters. Therefore, we experimented with this particular configuration of freezing the first three layers of the model. Building on this study, we also experimented with freezing the feature extractor for further reduction of parameters. The results presented in the Table 6 demonstrate that the outcomes of the freezed feature extractor are comparable to the unfreezed feature extractor. Thus, the rest of the experiments are done keeping the feature extractor layers frozen. Table 7 shows the comparison between the reduced learnable parameters.
@@
We have also compared our results from the top five frozen layers with the outcomes reported by Bayerl et al., 2023b . The Figure 4 illustrated that our model outcomes have outperformed the Wav2Vec2.0 proposed by Bayerl et al., 2023b with considerably less trainable parameters. In addition, it’s noteworthy that the Whisper model’s variant with 6 encoder layers was used in our experiments, and Wav2Vec2.0 model’s large variant which consists of 24 encoder layers was used for the respective task. Therefore, Wav2Vec2.0 large variant had 315.70 million trainable parameters, and our proposed whisper model configuration had 3.29 million trainable parameters.
@@
4.1 Theoretical and Practical Implications
This study validates the effectiveness of Whisper in multi-stuttered speech classification. The exploration of layer freezing strategies has revealed insights into the features that contribute most to stuttered speech classification. It has been identified through experimentation in our study that freezing the first five layers yields the most optimal results. The comparative analysis with the state-of-the-art Wav2Vec2.0 implies the superiority of Whisper in terms of both trainable parameters and F1-scores. In addition, the Whisper’s approach to acquire mel-spectrograms for feature extraction could contributing significantly contribute to better performance than Wav2Vec2.0, as it mimics human auditory system.
Since the proposed approach has outperformed Wav2Vec2.0 with significantly fewer trainable parameters, it suggests the potential for an automated assessment system to aid speech-language pathologists in diagnosis. The resource-efficient nature of this study enhances its feasibility for real-world applications. Researchers and developers can leverage these findings to improve results and expand into different languages.
@@ Acknowledgement
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this paper, the authors used Chat GPT 3.5 in order to rephrase less than 5% of the paper to improve its readability. After using this service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
[…]